# CPS-Guard: Framework for Dependability Assurance of AI- and LLM-Based Cyber-Physical Systems

### v1

@@ Abstract
Cyber-Physical Systems (CPS) increasingly depend on advanced AI techniques to operate in critical applications. However, traditional verification and validation methods often struggle to handle the unpredictable and dynamic nature of AI components. In this paper, we introduce CPS-Guard, a novel framework that employs multi-role orchestration to automate the iterative assurance process for AI-powered CPS. By assigning specialized roles (e.g., safety monitoring, security assessment, fault injection, and recovery planning) to dedicated agents within a simulated environment, CPS-Guard continuously evaluates and refines AI behavior against a range of dependability requirements. We demonstrate the framework through a case study involving an autonomous vehicle navigating an intersection with an AI-based planner. Our results show that CPS-Guard effectively detects vulnerabilities, manages performance impacts, and supports adaptive recovery strategies, thereby offering a structured and extensible solution for rigorous V&V in safety- and security-critical systems.
@@ I Introduction
To meet these needs, we present CPS-Guard, a novel framework centered around multi-role orchestration to iteratively assure the dependability of AI components within CPS. CPS-Guard assigns roles to various agents (AI models, algorithms, formal checkers, etc.) to work together within a simulated CPS environment. The agents continuously assess, challenge, and refine the performance of the primary AI component according to the defined dependability requirements.
@@
By orchestrating these roles in a coordinated manner, CPS-Guard aims to provide a more rigorous approach to V&V for AI in CPS than ad-hoc testing or isolated component analysis.
@@ II-C Runtime Assurance and Monitoring
Runtime Assurance (RA) techniques aim to ensure safety during operation by employment monitors and safety controllers [21, 22]. For example, the Simplex architecture [21] switches between a complex primary controller and a simpler, verified safety controller as needed. Runtime verification methods assess execution traces against formal specifications using temporal logic such as LTL, MTL, or STL [23, 24]. In contrast to these approaches, CPS-Guard utilizes multiple coordinated roles in a closed-loop assurance process. This goes beyond just switching to backup controllers.
@@ II-D Multi-Agent Systems (MAS) for Simulation and V&V
Multi-agent systems have been used to simulate complex systems and test control strategies in CPS [25]. Agent-based modeling has been used to mimic CPS behavior which focuses on autonomous driving scenarios. These approaches simply utilize agent interactions for simulation. On the other hand, CPS-Guard assigns specific V&V roles that target dependability assurance rather than general-purpose simulation or emergent behavior analysis.
@@ II-E LLM and AI Orchestration Frameworks
Recent frameworks such as LangChain [26] and LlamaIndex [27] have shown how to compose LLM calls with external tools and data to build applications. AutoGen [28] further supports complex workflows involving multiple LLM agents. Although these orchestration frameworks demonstrate considerable power, they are primarily designed for application development and general AI collaboration. Their capabilities for formal requirement checking, CPS simulation, and structured assurance loops are limited compared to the dedicated design of CPS-Guard. This often forces researchers to develop custom assurance loops and V&V steps from the ground up, which is highly time consuming. Earlier work, such as the LLMOrchestrator [29], provided a basic generator-verifier structure that CPS-Guard significantly expands by incorporating multiple V&V roles.
II-F Positioning CPS-Guard
CPS-Guard builds on ideas from the aforementioned fields while establishing a new niche. It adopts the orchestration paradigm from AI frameworks, re-purposing it specifically for dependability V&V. The frameworks employs a multi-role architecture inspired by multi-agent systems. It assigns roles that focus on safety monitoring, security testing, performance evaluation, fault injection, and recovery planning. Unlike approaches that rely solely on runtime monitoring and simple controller switching, CPS-Guard implements an iterative, closed-loop process that integrates tightly with CPS simulation environments as a robust solution for the complex challenges of V&V in AI-based CPS.
III The CPS-Guard Framework
CPS-Guard is a Python-based framework designed to structure and automate the iterative V&V process for AI components within simulated CPS environments. It employs a multi-role orchestration paradigm where specialized computational agents, termed Roles, collaborate to assess the behavior of the primary AI component Under Test (AUT) against dependability requirements.
@@ III-A Core Architecture
The CPS-Guard architecture, illustrated in Fig. 1, centers around an Orchestration Controller that manages the interaction between various Roles, an Environment Interface connecting to the CPS simulator, a State Manager maintaining shared state, and a Dependability Metrics tracker.
@@
Figure 1: Overview of the CPS-Guard architecture. Sensor data from the CPS is collected by the Environment Interface and organized by the State Manager. The Orchestration Controller coordinates specialized Roles that generate and refine actions through a dedicated Action Execution module. These actions are fed back into the system to complete a closed-loop assurance process, while Metrics are concurrently tracked for continuous verification and validation
@@ III-B Key Components
CPS-Guard consists of five key components.
@@ III-C Iterative Orchestration Workflow
The CPS-Guard workflow can be customized as needed. A typical execution cycle proceeds as follows:
@@ III-D Extensibility
CPS-Guard is designed for extensibility. Users can:
@@ IV Use Case: Autonomous Vehicles
To demonstrate the capabilities of CPS-Guard, we apply it to a challenging V&V scenario: ensuring the safe and secure navigation of an unsignalized urban intersection by an autonomous vehicle (AV) equipped with an AI-based planning module.
@@ IV-A Scenario Description
Figure 3: High-level orchestration of AI-based generator and other roles within CPS-Guard. The CARLA environment supplies navigation (map, traffic, waypoints) and sensor data. The SecurityAssessor can inject faults through the FaultInjector. These data streams, alongside the running state, feed into a prompt templater to generate a textual representation for the Llama 3.2 11B model. Camera views are passed directly to the LLM. Llama 3.2 generates both control outputs and corresponding explanations. The Action Execution module then applies these outputs to the environment. In parallel, the RecoveryPlanner employs geometric checks and determines whether or not to employ the emergency brake, which overrides all other actions. The running state is updated via the StateManager to include past actions and associated CoT explanations. The PerformanceOracle and SafetyMonitor track for performance.
@@ IV-A1 Rationale for an LLM-Based Planner
Existing AV planning stacks use domain-specific rule sets. While this is effective, it limits our ability to stress-test CPS-Guard across diverse AI use cases. We therefore deliberately use an LLM-based planner, which is relatively weaker in this area, to 1) show the framework can wrap any black-box decision module, 2) surface failure modes that traditional frameworks would not find, and 3) to demonstrate how DURA-CPS’s different roles interact when the AI’s internal logic is opaque.
IV-B CPS-Guard Configuration
We configured CPS-Guard with the following roles instantiated for this scenario (Figure 3):
@@ IV-B1 Integration
The CARLA simulator [31] was used via a custom CPS-Guard CarlaInterface (Figure 2). The StateManager tracked perceived object lists (from CARLA’s sensors), proposed actions, and associated CoT explanations.
@@ IV-D Expected Outcomes and Metrics
We used CPS-Guard to assess:
@@ V Results and Analysis
This section presents the results from executing the autonomous intersection navigation use case with CPS-Guard across the defined scenarios (15 runs per scenario, total 90 runs).
@@ V-A Safety Assessment
CPS-Guard effectively identified scenarios where the LLM planner proposed potentially unsafe actions.
@@ V-B Security Assessment and Resilience
CPS-Guard’s roles allowed systematic injection and observation of the AI’s vulnerability to sensor data manipulation.
@@ V-D Recovery Effectiveness
V-E CPS-Guard Contribution Analysis
This use case shows how CPS-Guard facilitates V&V:
@@ VI Discussion
The results from the autonomous intersection navigation use case provide insights into the challenges of assuring AI dependability in complex CPS and demonstrate the utility of the CPS-Guard framework’s multi-role orchestration approach.
@@ VI-A Implications for AI/LLM Assurance in CPS
Security attacks directly impacted safety (by inducing unsafe reactions or hesitation leading to secondary conflicts) and performance (by causing excessive caution or gridlock). This demonstrates the need for V&V frameworks like CPS-Guard that can assess these attributes concurrently and analyze their interactions, rather than treating them in isolation. The partial success of the simple recovery mechanism shows the importance of runtime assurance loops but also shows the need for more advanced monitoring and recovery strategies tailored to specific failure modes.
VI-B Effectiveness and Role of CPS-Guard
CPS-Guard proved effective in structuring this complex V&V task. Its key strengths observed in the use case include:
@@ VI-B Effectiveness and Role of CPS-Guard
CPS-Guard acts as an "in-the-loop V&V orchestrator," allowing engineers to configure a virtual team of specialized agents that continuously probe and assess the AI AUT within its simulated operational context. This approach facilitates identifying weaknesses, understanding interactions between dependability facets, and evaluating the effectiveness of assurance mechanisms like monitors and recovery planners.
@@ VI-C Limitations and Future Work
  • Sim-to-Real Gap: As with all simulation-based V&V, transferring findings to the real world requires caution. Validating CPS-Guard and role implementations on physical platforms or high-fidelity digital twins is crucial.
@@
    Integrating CPS-Guard with hardware-in-the-loop (HIL) setups to evaluate its effectiveness.
@@
    Applying CPS-Guard to other domains, including industrial robotics and dependable agricultural automation (e.g., safe human-robot collaboration in smart farming).
@@
    Investigating optimizations and removing bottlenecks that limit CPS-Guard’s effectiveness in out-of-sim tests.
@@
    Developing specialized assessment metrics tailored to LLM-specific failure modes, such as hallucination, and integrating these into the CPS-Guard framework to improve LLM-based component assurance.
@@ VII Conclusion
This paper introduces CPS-Guard, a framework leveraging multi-role orchestration to structure and automate the iterative assurance process for AI-based CPS. By assigning specialized V&V functions to roles within a simulated CPS environment, CPS-Guard systematically evaluates AI behavior against a range of dependability requirements. Its iterative and closed-loop approach allows analysis of complex interactions and continuous evaluation of runtime assurance mechanisms.
We demonstrated the utility of CPS-Guard in a case study on autonomous vehicle intersection navigation with an LLM-based planner. The case study shows how CPS-Guard can effectively inject faults, detect safety and security vulnerabilities, assess performance impacts, and evaluate recovery actions. CPS-Guard provides quantitative metrics for assurance.
In conclusion, CPS-Guard offers a practical method for the V&V of AI in CPS where safety and security are critical. Further work should focus on the identified challenges like specification effort and bridging the simulation-to-reality gap.

### v2

# DURA-CPS: A Multi-Role Orchestrator for Dependability Assurance in LLM-Enabled Cyber-Physical Systems

@@ Abstract
Cyber-Physical Systems (CPS) increasingly depend on advanced AI techniques to operate in critical applications. However, traditional verification and validation methods often struggle to handle the unpredictable and dynamic nature of AI components. In this paper, we introduce DURA-CPS, a novel framework that employs multi-role orchestration to automate the iterative assurance process for AI-powered CPS. By assigning specialized roles (e.g., safety monitoring, security assessment, fault injection, and recovery planning) to dedicated agents within a simulated environment, DURA-CPS continuously evaluates and refines AI behavior against a range of dependability requirements. We demonstrate the framework through a case study involving an autonomous vehicle navigating an intersection with an AI-based planner. Our results show that DURA-CPS effectively detects vulnerabilities, manages performance impacts, and supports adaptive recovery strategies, thereby offering a structured and extensible solution for rigorous V&V in safety- and security-critical systems.
@@ I Introduction
To meet these needs, we present DURA-CPS, a novel framework centered around multi-role orchestration to iteratively assure the dependability of AI components within CPS. DURA-CPS assigns roles to various agents (AI models, algorithms, formal checkers, etc.) to work together within a simulated CPS environment. The agents continuously assess, challenge, and refine the performance of the primary AI component according to the defined dependability requirements.
@@
By orchestrating these roles in a coordinated manner, DURA-CPS aims to provide a more rigorous approach to V&V for AI in CPS than ad-hoc testing or isolated component analysis.
@@ II-C Runtime Assurance and Monitoring
Runtime Assurance (RA) techniques aim to ensure safety during operation by employment monitors and safety controllers [21, 22]. For example, the Simplex architecture [21] switches between a complex primary controller and a simpler, verified safety controller as needed. Runtime verification methods assess execution traces against formal specifications using temporal logic such as LTL, MTL, or STL [23, 24]. In contrast to these approaches, DURA-CPS utilizes multiple coordinated roles in a closed-loop assurance process. This goes beyond just switching to backup controllers.
@@ II-D Multi-Agent Systems (MAS) for Simulation and V&V
Multi-agent systems have been used to simulate complex systems and test control strategies in CPS [25]. Agent-based modeling has been used to mimic CPS behavior which focuses on autonomous driving scenarios. These approaches simply utilize agent interactions for simulation. On the other hand, DURA-CPS assigns specific V&V roles that target dependability assurance rather than general-purpose simulation or emergent behavior analysis.
@@ II-E LLM and AI Orchestration Frameworks
Recent frameworks such as LangChain [26] and LlamaIndex [27] have shown how to compose LLM calls with external tools and data to build applications. AutoGen [28] further supports complex workflows involving multiple LLM agents. Although these orchestration frameworks demonstrate considerable power, they are primarily designed for application development and general AI collaboration. Their capabilities for formal requirement checking, CPS simulation, and structured assurance loops are limited compared to the dedicated design of DURA-CPS. This often forces researchers to develop custom assurance loops and V&V steps from the ground up, which is highly time consuming. Earlier work, such as the LLMOrchestrator [29], provided a basic generator-verifier structure that DURA-CPS significantly expands by incorporating multiple V&V roles.
II-F Positioning DURA-CPS
DURA-CPS builds on ideas from the aforementioned fields while establishing a new niche. It adopts the orchestration paradigm from AI frameworks, re-purposing it specifically for dependability V&V. The frameworks employs a multi-role architecture inspired by multi-agent systems. It assigns roles that focus on safety monitoring, security testing, performance evaluation, fault injection, and recovery planning. Unlike approaches that rely solely on runtime monitoring and simple controller switching, DURA-CPS implements an iterative, closed-loop process that integrates tightly with CPS simulation environments as a robust solution for the complex challenges of V&V in AI-based CPS.
III The DURA-CPS Framework
DURA-CPS is a Python-based framework designed to structure and automate the iterative V&V process for AI components within simulated CPS environments. It employs a multi-role orchestration paradigm where specialized computational agents, termed Roles, collaborate to assess the behavior of the primary AI component Under Test (AUT) against dependability requirements.
@@ III-A Core Architecture
The DURA-CPS architecture, illustrated in Fig. 1, centers around an Orchestration Controller that manages the interaction between various Roles, an Environment Interface connecting to the CPS simulator, a State Manager maintaining shared state, and a Dependability Metrics tracker.
@@
Figure 1: Overview of the DURA-CPS architecture. Sensor data from the CPS is collected by the Environment Interface and organized by the State Manager. The Orchestration Controller coordinates specialized Roles that generate and refine actions through a dedicated Action Execution module. These actions are fed back into the system to complete a closed-loop assurance process, while Metrics are concurrently tracked for continuous verification and validation
@@ III-B Key Components
DURA-CPS consists of five key components.
@@ III-C Iterative Orchestration Workflow
The DURA-CPS workflow can be customized as needed. A typical execution cycle proceeds as follows:
@@ III-D Extensibility
DURA-CPS is designed for extensibility. Users can:
@@ IV Use Case: Autonomous Vehicles
To demonstrate the capabilities of DURA-CPS, we apply it to a challenging V&V scenario: ensuring the safe and secure navigation of an unsignalized urban intersection by an autonomous vehicle (AV) equipped with an AI-based planning module.
@@ IV-A Scenario Description
Figure 3: High-level orchestration of AI-based generator and other roles within DURA-CPS. The CARLA environment supplies navigation (map, traffic, waypoints) and sensor data. The SecurityAssessor can inject faults through the FaultInjector. These data streams, alongside the running state, feed into a prompt templater to generate a textual representation for the Llama 3.2 11B model. Camera views are passed directly to the LLM. Llama 3.2 generates both control outputs and corresponding explanations. The Action Execution module then applies these outputs to the environment. In parallel, the RecoveryPlanner employs geometric checks and determines whether or not to employ the emergency brake, which overrides all other actions. The running state is updated via the StateManager to include past actions and associated CoT explanations. The PerformanceOracle and SafetyMonitor track for performance.
@@ IV-A1 Rationale for an LLM-Based Planner
Existing AV planning stacks use domain-specific rule sets. While this is effective, it limits our ability to stress-test DURA-CPS across diverse AI use cases. We therefore deliberately use an LLM-based planner, which is relatively weaker in this area, to 1) show the framework can wrap any black-box decision module, 2) surface failure modes that traditional frameworks would not find, and 3) to demonstrate how DURA-CPS’s different roles interact when the AI’s internal logic is opaque.
IV-B DURA-CPS Configuration
We configured DURA-CPS with the following roles instantiated for this scenario (Figure 3):
@@ IV-B1 Integration
The CARLA simulator [31] was used via a custom DURA-CPS CarlaInterface (Figure 2). The StateManager tracked perceived object lists (from CARLA’s sensors), proposed actions, and associated CoT explanations.
@@ IV-D Expected Outcomes and Metrics
We used DURA-CPS to assess:
@@ V Results and Analysis
This section presents the results from executing the autonomous intersection navigation use case with DURA-CPS across the defined scenarios (15 runs per scenario, total 90 runs).
@@ V-A Safety Assessment
DURA-CPS effectively identified scenarios where the LLM planner proposed potentially unsafe actions.
@@ V-B Security Assessment and Resilience
DURA-CPS’s roles allowed systematic injection and observation of the AI’s vulnerability to sensor data manipulation.
@@ V-D Recovery Effectiveness
V-E DURA-CPS Contribution Analysis
This use case shows how DURA-CPS facilitates V&V:
@@ VI Discussion
The results from the autonomous intersection navigation use case provide insights into the challenges of assuring AI dependability in complex CPS and demonstrate the utility of the DURA-CPS framework’s multi-role orchestration approach.
@@ VI-A Implications for AI/LLM Assurance in CPS
Security attacks directly impacted safety (by inducing unsafe reactions or hesitation leading to secondary conflicts) and performance (by causing excessive caution or gridlock). This demonstrates the need for V&V frameworks like DURA-CPS that can assess these attributes concurrently and analyze their interactions, rather than treating them in isolation. The partial success of the simple recovery mechanism shows the importance of runtime assurance loops but also shows the need for more advanced monitoring and recovery strategies tailored to specific failure modes.
VI-B Effectiveness and Role of DURA-CPS
DURA-CPS proved effective in structuring this complex V&V task. Its key strengths observed in the use case include:
@@ VI-B Effectiveness and Role of DURA-CPS
DURA-CPS acts as an "in-the-loop V&V orchestrator," allowing engineers to configure a virtual team of specialized agents that continuously probe and assess the AI AUT within its simulated operational context. This approach facilitates identifying weaknesses, understanding interactions between dependability facets, and evaluating the effectiveness of assurance mechanisms like monitors and recovery planners.
@@ VI-C Limitations and Future Work
  • Sim-to-Real Gap: As with all simulation-based V&V, transferring findings to the real world requires caution. Validating DURA-CPS and role implementations on physical platforms or high-fidelity digital twins is crucial.
@@
    Integrating DURA-CPS with hardware-in-the-loop (HIL) setups to evaluate its effectiveness.
@@
    Applying DURA-CPS to other domains, including industrial robotics and dependable agricultural automation (e.g., safe human-robot collaboration in smart farming).
@@
    Investigating optimizations and removing bottlenecks that limit DURA-CPS’s effectiveness in out-of-sim tests.
@@
    Developing specialized assessment metrics tailored to LLM-specific failure modes, such as hallucination, and integrating these into the DURA-CPS framework to improve LLM-based component assurance.
@@ VII Conclusion
This paper introduces DURA-CPS, a framework leveraging multi-role orchestration to structure and automate the iterative assurance process for AI-based CPS. By assigning specialized V&V functions to roles within a simulated CPS environment, DURA-CPS systematically evaluates AI behavior against a range of dependability requirements. Its iterative and closed-loop approach allows analysis of complex interactions and continuous evaluation of runtime assurance mechanisms.
We demonstrated the utility of DURA-CPS in a case study on autonomous vehicle intersection navigation with an LLM-based planner. The case study shows how DURA-CPS can effectively inject faults, detect safety and security vulnerabilities, assess performance impacts, and evaluate recovery actions. DURA-CPS provides quantitative metrics for assurance.
In conclusion, DURA-CPS offers a practical method for the V&V of AI in CPS where safety and security are critical. Further work should focus on the identified challenges like specification effort and bridging the simulation-to-reality gap.